Graph the function: g(x)=x^2-7x+10.
Find the axis of symmetry, the vertex, the max and min, and the range.
My teacher gave me this problem for homework. I don"t know if he got it from a text book.

Answer by venugopalramana(3286)

Graph the function: g(x)=x^2-7x+10.
Find the axis of symmetry, the vertex, the max and min, and the range.
G(X) = [X^2-2*(X)*(7/2)+(7/2)^2]-(7/2)^2+10
= [X-(7/2)]^2+[10-49/4]=[X-3.5]^2-9/4
HENCE AXIS OF SYMMETRY IS X-3.5=0...OR X=3.5
VERTEX IS AT X=3.5...AND Y=-9/4...SO VERTEX IS (3.5,-9/4)
SINCE (X-3.5)^2 IS ALWAYS +VE,THE FUNCTION HAS A MINIMUM AT THE VERTEX ..ITS VALUE
BEING -9/4
MAXIMUM IS INFINITY.
RANGE IS FROM -9/4 TO INFINITY.(-9/4,INFINITY)








My teacher gave me this problem for homework. I don"t know if he got it from a text book.